OpenStudy (anonymous):

Choose the equation of the line passing through the point (1, -5) and perpendicular to y = x 1/5 - 10. y = 5x + 10 y = 5x y = -5x - 10 y = -5x

OpenStudy (campbell_st):

gradient of the line is \[m _{l} = \frac{1}{5}\] for a perpendicular \[m _{p}=-\frac{1}{m _{l}}\] to the gradient of the perpendicular is \[m _{p}= -5\] in point slope form the line is \[y -(-5) = -5(x - 1)\] simplify to find the equation of your line

OpenStudy (anonymous):

Is it y= -5x-10?

OpenStudy (campbell_st):

1st simplifcation is y + 5 = -5x + 5

OpenStudy (anonymous):

y=-5x!

OpenStudy (campbell_st):

correct